# Self-Supervised Models in Automatic Whispered Speech Recognition

### v1

@@ Abstract
In automatic speech recognition, any factor that alters the acoustic properties of speech can pose a challenge to the system’s performance. This paper presents a novel approach for automatic whispered speech recognition in the Irish dialect using the self-supervised WavLM model. Conventional automatic speech recognition systems often fail to accurately recognise whispered speech due to its distinct acoustic properties and the scarcity of relevant training data. To address this challenge, we utilized a pre-trained WavLM model, fine-tuned with a combination of whispered and normal speech data from the wTIMIT and CHAINS datasets, which include the English language in Singaporean and Irish dialects, respectively. Our baseline evaluation with the OpenAI Whisper model highlighted its limitations, achieving a Word Error Rate (WER) of 18.8% on whispered speech. In contrast, the proposed WavLM-based system significantly improved performance, achieving a WER of 9.22%. These results demonstrate the efficacy of our approach in recognising whispered speech and underscore the importance of tailored acoustic modeling for robust automatic speech recognition systems. This study provides valuable insights into developing effective automatic speech recognition solutions for challenging speech affected by whisper and dialect. The source codes for this paper are freely available  [^1] GitHub Repository for Whisper transcription. [^2] GitHub Repository for WavLM fine-tuning..
@@ I Introduction
Recently, self-supervised models have been extensively utilized in various speech processing fields, including speech recognition, speaker recognition, and speaker diarization [1]. These models have notably enhanced the efficiency of such systems, particularly for the English language. The training of these models typically involves large datasets comprising speech from different languages, predominantly normal speech. However, these models exhibit diminished performance when processing abnormal speech types such as whispered, reverberant, and disordered speech, as these variants are not represented in the training data.
Whispered speech is a distinct mode of speech production that most individuals can produce naturally and at their discretion. It is often used for privacy or to avoid disturbing others in environments like libraries. Additionally, whispered speech may be a compulsion for individuals with laryngeal or vocal cord damage. The acoustic properties of whispered speech differ substantially from those of normal speech, presenting significant challenges for conventional speech processing systems.
Furthermore, the presence of different dialects within a language can adversely impact the performance of automatic speech recognition (ASR) systems trained on standard dialects. Variations in dialect result in significant changes in the acoustic parameters of speech, thereby reducing the accuracy of ASR systems when encountering non-standard dialectal speech.
In this work, we aim to design an ASR system for whispered speech in the Irish dialect using the powerful self-supervised model, WavLM. Among various self-supervised models, WavLM has demonstrated superior performance in different scenarios, particularly in speaker recognition, where the acoustic characteristics of the speaker are more critical than the speech content. Consequently, WavLM has been selected to address the challenge of recognising whispered speech in the Irish dialect.
To illustrate the limitations of conventional speech recognition systems, we will also employ the Whisper model introduced by OpenAI [2], a widely acclaimed model in normal speech recognition. In this article, the Whisper model developed by OpenAI is italicized to differentiate it easily from the concept of whispered speech.
Developing a speech recognition system necessitates a robust and comprehensive dataset. However, research on whispered speech is limited primarily due to the scarcity of such datasets. Currently, only a few whispered speech datasets, such as wTIMIT [3] and CHAINS [4], are available and primarily cater to the English language. While some processes can effectively address some challenges, such as producing reverberant speech data [5] as a data augmentation method, they have not yet proven suitable for whispered speech conversion. Despite some efforts in the literature to develop methods for automatically converting normal speech to whispered speech, these artificially generated whispered speech samples tend to be unnatural and noisy [6].
To mitigate the impact of data scarcity, fine-tuning self-supervised models presents a viable solution compared to end-to-end (E2E) methods based on conformers and transformers [7], as suggested in several studies. This approach leverages the existing knowledge encoded in pre-trained models to enhance performance with limited training data.
@@
Previous work has explored various methods for recognising whispered speech, considering its unique characteristics and the challenge of limited data availability. As one of the first works in whispered speech recognition, [8], designed a system based on hidden Markov model and mel frequency cepstral coefficient features. Some works were also focused on improving the features [9] and adapting the acoustic model [10, 11, 12]. However, Several works focused on improving pre-processing [13, 14] to enhance system performance for whispered speech.
@@ III-A WavLM
The WavLM network, a self-supervised learning (SSL) model introduced by Microsoft in 2021, aims to provide a comprehensive solution for various speech processing applications. WavLM was designed and trained for full-stack speech processing differently from other SSL models, achieving superior performance in tasks such as speaker recognition and separation, which demand an efficient understanding of acoustic models. The training process of this model involved masked speech prediction and denoising, which were performed jointly. This dual-task training strategy enables the model to predict masked speech content while learning to denoise, thereby equipping it with the ability to handle multiple speech processing scenarios beyond ASR.
@@
III-B Whisper Dataset
The CHAINS dataset is utilized to evaluate the proposed system. This dataset includes both normal speech, referred to as solo, and whispered speech, known as whsp. The dataset comprises recordings from 36 speakers, dominantly with Irish dialects. To develop a model capable of accurately capturing the acoustic features of the Irish dialect, the normal speech portion of this dataset is used for system training, while the whispered speech portion is employed for evaluation. Additionally, to improve the model’s performance on a whispered speech by learning its unique acoustic characteristics, the whispered speech portion of the wTIMIT dataset, which includes Singaporean and US dialects, is also used as auxiliary data in the training process.
In this evaluation, only the Singaporean dialect whispered speech from wTIMIT is used for training, ensuring that the system encounters different dialects from those used during pre-training, thereby fostering a more general acoustic model. Each speaker’s evaluation data involves 33 whispered utterances from the CHAINS dataset. Further details of these datasets are presented in Table I.
@@ IV Experimental Results
The ASR results using the Large-v2 version of the Whisper model as the baseline system are presented in Table II. This system achieved a WER of 16.69% for normal and 18.8% for whispered speech parts of the CHAINS dataset. The speech texts in the dataset are identical for both speech styles. As shown, the system experienced a drop in performance when dealing with whispered speech. In addition, the performance of the Irish dialect was lower than that of the typical English [2]. These results indicate that whispered speech with a different dialect presents substantial modeling complexities for a system trained predominantly on normal speech data, underscoring the need for a dedicated system for whispered speech recognition.
@@
TABLE II: WERof pre-trained Whisper model for normal and whispered speech style
Pre-trained Model Normal speech Whispered speech
OpenAI Whisper 16.69% 18.80%
TABLE III: WERof fine-tuned WavLM model for whispered speech style with two decoding methods
Fine-tuned Model Greedy search Beam search
WavLM 9.28% 9.22%
The proposed WavLM-based system achieved a WER of 9.28% using the greedy search algorithm and 9.22% based on the beam search algorithm. These results demonstrate that the system effectively learned the acoustic parameters of whispered speech, significantly outperforming the OpenAI Whisper model baseline system. This improvement can be attributed to including whispered speech data in the training phase. Additionally, the Irish dialect, which posed a challenge for the baseline system, was effectively addressed during training using normal speech data from the CHAINS dataset. The choice of the WavLM network proved advantageous due to its suitability for training on small datasets and its superior acoustic modeling capabilities.
@@ V Conclusion
In this article, we proposed a system for speech recognition under the challenging conditions of whispered speech and the Irish dialect. To address these challenges, we utilized training data with an acoustic space similar to the test conditions, enabling the system to effectively model the complex acoustic characteristics in the test data. The results demonstrate that the proposed system, with a WER of 9.22%, outperforms the powerful OpenAI Whisper system, which served as a baseline. Our findings indicate that the proposed system achieves superior performance on the CHAINS dataset for whispered speech recognition.

### v2

# Leveraging Self-Supervised Models for Automatic Whispered Speech Recognition

@@ Abstract
In automatic speech recognition, any factor that alters the acoustic properties of speech can pose a challenge to the system’s performance. This paper presents a novel approach for automatic whispered speech recognition in the Irish dialect using the self-supervised WavLM model. Conventional automatic speech recognition systems often fail to accurately recognise whispered speech due to its distinct acoustic properties and the scarcity of relevant training data. To address this challenge, we utilized a pre-trained WavLM model, fine-tuned with a combination of whispered and normal speech data from the wTIMIT and CHAINS datasets, which include the English language in Singaporean and Irish dialects, respectively. Our baseline evaluation with the OpenAI Whisper model highlighted its limitations, achieving a Word Error Rate (WER) of 18.8% and a Character Error Rate (CER) of 4.24% on whispered speech. In contrast, the proposed WavLM-based system significantly improved performance, achieving a WER of 9.22% and a CER of 2.59%. These results demonstrate the efficacy of our approach in recognising whispered speech and underscore the importance of tailored acoustic modeling for robust automatic speech recognition systems. This study provides valuable insights into developing effective automatic speech recognition solutions for challenging speech affected by whisper and dialect. The source codes for this paper are freely available  [^1] GitHub Repository for Whisper transcription. [^2] GitHub Repository for WavLM fine-tuning..
Keywords— Whispered Speech Recognition, self-supervised models, Whisper model, WavLM.
@@ I Introduction
Recent advancements in self-supervised models have significantly impacted various speech processing domains, such as speech recognition, speaker recognition [1, 2], and speaker diarization [3]. These models, especially in English, have substantially improved system efficiency by leveraging large multilingual datasets, primarily featuring normal speech. However, these models struggle when processing atypical forms of speech, including whispered, reverberant [4], and disordered speech [5, 6], as such variants are underrepresented in the training data.
Whispered speech, a distinct mode of communication, is naturally produced by most individuals for privacy or when vocal cords are compromised. Its acoustic characteristics differ markedly from normal speech, which poses a unique challenge for traditional speech recognition systems. These challenges highlight the importance of developing specialized automatic speech recognition (ASR) systems to handle whispered speech, especially considering its practical applications in healthcare and privacy-sensitive environments.
Dialectal variations, in addition to whispered speech, further complicate speech recognition. Dialects introduce significant changes in acoustic parameters, which reduce the accuracy of ASR systems trained predominantly on standard dialects. In particular, whispered speech in non-standard dialects, such as the Irish dialect, presents an even more significant challenge. No multi-modal dataset exists for whispered speech in any dialect, further compounding the problem.
This research addresses these critical gaps by focusing on the original and novel goal of developing an ASR system specifically designed to recognize whispered speech in the Irish dialect. Leveraging the capabilities of the self-supervised WavLM model, which has shown superior performance across various tasks, this study aims to enhance the system’s ability to identify whispered speech accurately. By incorporating dialectal variations into the model’s training, this research aims to solve a largely underexplored area.
To further emphasize the limitations of traditional models, we will utilize OpenAI’s Whisper model [7], a state-of-the-art system renowned for its performance in standard speech recognition tasks. Comparing Whisper to WavLM in the context of whispered speech highlights the current limitations and underscores the necessity of specialized models for whispered and dialectal speech.
The development of an effective ASR system requires a comprehensive and diverse dataset. However, research on whispered speech remains limited due to a scarcity of high-quality data. Existing datasets, such as wTIMIT [8] and CHAINS [9], primarily focus on English and fail to address the nuances of whispered speech in other languages or dialects. Furthermore, while data augmentation techniques like generating reverberant speech [10] have proven successful for normal speech, they are inadequate for whispered speech conversion. Efforts to artificially convert normal speech to whispered speech remain problematic due to the unnatural and noisy outputs [11].
Given these challenges, fine-tuning pre-trained self-supervised models offers a practical solution to the scarcity of whispered speech data. This method, as supported by recent research [12], allows models to leverage existing knowledge and adapt to new tasks, offering a more viable alternative than traditional end-to-end (E2E) systems based on conformers or transformers. In this study, we aim to address the data scarcity challenge and demonstrate the effectiveness of fine-tuning techniques in developing robust ASR systems for whispered speech.
@@
Previous work has explored various methods for recognising whispered speech, considering its unique characteristics and the challenge of limited data availability. As one of the first works in whispered speech recognition, [13], designed a system based on hidden Markov model and mel frequency cepstral coefficient features. Some works were also focused on improving the features [14, 15] and adapting the acoustic model [16, 17, 18]. However, Several works focused on improving pre-processing [19, 20] to enhance system performance for whispered speech.
@@ III-A WavLM
The WavLM model [40], a self-supervised learning (SSL) model introduced by Microsoft in 2021, aims to provide a comprehensive solution for various speech processing applications. WavLM was designed and trained for full-stack speech processing differently from other SSL models, achieving superior performance in tasks such as speaker recognition and separation, which demand an efficient understanding of acoustic models. The training process of this model involved masked speech prediction and denoising, which were performed jointly. This dual-task training strategy enables the model to predict masked speech content while learning to denoise, thereby equipping it with the ability to handle multiple speech processing scenarios beyond ASR.
@@
III-B Whispered Speech Dataset
The CHAINS dataset [9] is utilized to evaluate the proposed system. This dataset includes both normal speech, referred to as solo, and whispered speech, known as whsp. The dataset comprises recordings from 36 speakers, dominantly with Irish dialects. To develop a model capable of accurately capturing the acoustic features of the Irish dialect, the normal speech portion of this dataset is used for system training, while the whispered speech portion is employed for evaluation. Additionally, to improve the model’s performance on a whispered speech by learning its unique acoustic characteristics, the whispered and normal speech portion of the wTIMIT dataset [8], which is in Singaporean dialect, is also used as auxiliary data in the training process.
In this evaluation, only the Singaporean dialect speech from wTIMIT is used for training, ensuring that the system encounters different dialects from those used during pre-training, thereby fostering a more general acoustic model. Each speaker’s evaluation data involves 33 whispered utterances from the CHAINS dataset. Further details of these datasets are presented in Table I.
@@ IV Experimental Results
The ASR results using the Large-v2 version of the Whisper model as the baseline system are presented in Table II. This system achieved a WER of 16.69% and CER of 3.81% for normal and WER of 18.8% and CER of 4.24% for whispered speech parts of the CHAINS dataset. The speech texts in the dataset are identical for both speech styles. As shown, the system experienced a drop in performance when dealing with whispered speech. In addition, the performance of the Irish dialect was lower than that of the typical English [7]. These results indicate that whispered speech with a different dialect presents substantial modeling complexities for a system trained predominantly on normal speech data, underscoring the need for a dedicated system for whispered speech recognition.
@@
TABLE II: WERand CERof pre-trained Whisper model for normal and whispered speech styles
Metric Normal Speech Whispered Speech
WER16.69 18.80
CER3.81 4.24
TABLE III: WERand CERof fine-tuned WavLM model for whispered speech style with two decoding methods
Metric Greedy Search Beam Search
WER9.28 9.22
CER2.60 2.59
The proposed WavLM-based system achieved a WER of 9.28% using the greedy search algorithm and 9.22% based on the beam search algorithm, and the best achievement for the CER metric was 2.59% for greedy search. These results demonstrate that the system effectively learned the acoustic parameters of whispered speech, significantly outperforming the OpenAI Whisper model baseline system. This improvement can be attributed to including whispered speech data in the training phase. Additionally, the Irish dialect, which posed a challenge for the baseline system, was effectively addressed during training using normal speech data from the CHAINS dataset. The choice of the WavLM network proved advantageous due to its suitability for training on small datasets and its superior acoustic modeling capabilities.
@@ V Conclusion
In this article, we proposed a speech recognition system to handle the challenging conditions of whispered speech and the Irish dialect. By using training data that closely aligns with the acoustic characteristics of the test conditions, our system effectively modeled the complex nature of whispered speech. The results demonstrate the superiority of our system, achieving a Word Error Rate (WER) of 9.22%, significantly outperforming the OpenAI Whisper system, which served as our baseline. Based on our knowledge, this is the first work on whispered speech recognition on the CHAINS dataset. Furthermore, our findings show that the system delivers strong performance on the CHAINS dataset, particularly in the context of whispered speech recognition.
However, several limitations must be acknowledged. The primary constraint of this research is the scarcity of whispered speech datasets, especially those that include dialect variations. Although existing datasets like wTIMIT and CHAINS were used, they primarily focus on the English language and lack the diversity needed for robust, generalized performance across different dialects and speech conditions.
For future work, two key areas could be explored to enhance the performance and applicability of whispered speech recognition systems. First, utilizing other self-supervised models could offer insights into how different architectures handle whispered speech, potentially leading to further improvements in performance. Second, designing and developing a new, comprehensive dataset for whispered speech, particularly one that includes various dialects and languages, would be a critical step toward overcoming the dataset scarcity issue.